Grade 12Engineering Entrance Exams

A rocket with a lift-off mass 3.5 *10 to the4 kg is blasted upwards with an initial acceleration of 10 m/s2. Then the initial thrust of the blast is ?

sudhanshu
12 Years agoGrade 12

3 Answers

Psv Dan
9 Years ago
since it is initial thrust 
 
initial thrust= mass * acceleration (just a formula)
                        3.5 *104 x 10
=3.5*105
Santosh Narayanan
7 Years ago
F= m(g+a)                    m= 3.5 * 10kg
= 3.5 * 10(10 + 10)   g= 10 m/s2
=3.5 * 10* 20             a= 10 m/s2
=7.0 * 105                          Thrust=?
anushka
7 Years ago
here thrust force is responsible to accelerate the rocket so  initial thrust of the blast
=ma= 3.5×10^4×10
 
 

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago